OpenStudy (anonymous):

can sum1 plz help me.. find2 positive real numbers whose sum is 13 and whose product is a maximum! plzz n thnk u

Parth (parthkohli):

6 and 7?

OpenStudy (anonymous):

no its not :(

Parth (parthkohli):

Okay.. they can be decimals too... I get it.

Parth (parthkohli):

6.5 * 6.5... This one... Just in the middle

OpenStudy (anonymous):

yes thst it bt can u please do it step by step? nn wow tht was fast

OpenStudy (anonymous):

xy = 13 and S=x+y y=13/x -> S=x+13/x then Differentiate for max -> 6.5

Parth (parthkohli):

Yeah.. I noticed a pattern. If you take 12 and 1, you have the smallest, 11 and 2 gets a little bit bigger, 10 and 3 gets a little bit more bigger ....... 7 and 6, then 6.5 * 6.5 is the biggest one. This is really simple right?

OpenStudy (anonymous):

A.M > G.M 13/2 = 6.5

OpenStudy (anonymous):

Yes, it's symmetrical so min max same...

OpenStudy (anonymous):

thnk u guyz nn yes tht was simple bt i need to do it in steps showing my work.. n they want me to break it down and show it

Parth (parthkohli):

@khadija than an 8th grader won't help :P

OpenStudy (anonymous):

u r in 8 grade?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

The best way is to use the identity airthmetic mean >= geometric mean. and the rest is just one division.

OpenStudy (anonymous):

oh its ok then ;) n can sum then fo it using the quadratic equation

OpenStudy (anonymous):

can anybody do this question or nt?